OpenStudy (anonymous):

you have27 wooden cubes that he wants to paint to give to his little sister to use for blocks, but he needs to know the surface areas of the cubes so that he can figure out how much paint to buy. Each of the cubes has sides of 7 centimeters in length. What is the total surface area of all of the wooden cubes? Thanks :D

OpenStudy (lollylau):

surface area of one cube = 7*7*6 final ans. = 7 * 7 * 6 * 27.

OpenStudy (lollylau):

buy 1 full tin.

OpenStudy (lollylau):

just dip them in and its done

OpenStudy (anonymous):

one full tin what?

OpenStudy (anonymous):

one full tin what?

OpenStudy (anonymous):

one full tin what?

OpenStudy (lollylau):

paint, of course.